# PediaBench: A Comprehensive Chinese Pediatric Dataset for Benchmarking Large Language Models

### v2

@@ Abstract
The emergence of Large Language Models (LLMs) in the medical domain has stressed a compelling need for standard datasets to evaluate their question-answering (QA) performance. Although there have been several benchmark datasets for medical QA, they either cover common knowledge across different departments or are specific to another department rather than pediatrics. Moreover, some of them are limited to objective questions and do not measure the generation capacity of LLMs. Therefore, they cannot comprehensively assess the QA ability of LLMs in pediatrics. To fill this gap, we construct PediaBench, the first Chinese pediatric dataset for LLM evaluation. Specifically, it contains 4,565 objective questions and 1,632 subjective questions spanning 12 pediatric disease groups. It adopts an integrated scoring criterion based on different difficulty levels to thoroughly assess the proficiency of an LLM in instruction following, knowledge understanding, clinical case analysis, etc. Finally, we validate the effectiveness of PediaBench with extensive experiments on 20 open-source and commercial LLMs. Through an in-depth analysis of experimental results, we offer insights into the ability of LLMs to answer pediatric questions in the Chinese context, highlighting their limitations for further improvements. Our code and data are published at https://github.com/ACMISLab/PediaBench.
@@ 1 Introduction
To address the above challenges, in this paper, we introduce PediaBench, the first comprehensive Chinese benchmark dataset in pediatrics. As shown in Figure 1, PediaBench contains 4,565 objective questions and 1,632 subjective questions collected from various sources, such as the Chinese National Medical Licensing Examination (CNMLE), final exams of universities in medicine, pediatric disease diagnosis and treatment standards, and clinical guidelines. PediaBench encompasses five distinct types of questions, namely true or false, multiple choice, pairing, essay/short answer, and case analysis, across 12 typical pediatric disease groups. Furthermore, to provide an accurate evaluation of the performance of each LLM for QA in pediatrics, we use a scoring criterion that combines difficulty levels and automatic scoring. On the one hand, we assign different scores to objective questions based on the difficulty factor; on the other hand, we adopt the Language-Model-as-an-Examiner (LME) framework Bai et al. (2023b) to score the objective questions based on GPT-4o. Finally, we conduct extensive experiments with 20 LLMs on the PediaBench dataset and provide an in-depth analysis of the performance of LLMs for QA in pediatrics. Our main contributions are summarized as follows:
@@ 3.4 Evaluation Criteria
For ToF and MC, we use accuracy as a basic performance measure for each question. Then, we design a scoring scheme based on difficulty coefficient to assess the overall performance of each LLM. We calculate the difficulty coefficient $Dc_{i}$ of each question $i$ based on the accuracy of each LLM’s answer. Specifically, we have $Dc_{i}=1-\frac{\text{num}^{\text{acc}}_{i}}{\text{num}^{\text{llm}}_{i}}$, where $\text{num}^{\text{acc}}_{i}$ is the number of LLMs that correctly answer question $i$ and $\text{num}^{\text{llm}}_{i}$ is the total number of LLMs in the evaluation. All questions are then divided into four difficulty levels, namely simple, normal, difficult, and extreme, with different weights $w_{i}$ based on $Dc_{i}$ in the scoring scheme as follows:
@@
Figure 3 presents the number of questions at each difficulty level in ToF and MC. For PA, we assign a weight of $w_{i}=3$ to each question $i$ and use the following scoring rules: (1) A completely correct answer gets $3$ points; (2) a partially correct answer gets $1$ point; and (3) a completely incorrect answer does not get any points. Finally, the overall score of an LLM for ToF, MC, or PA is the weighted sum of the scores of all correctly answered questions.
@@
ES and CA are open questions with no unique answers. CA is even harder than ES because it may require the integration of multiple pieces of knowledge to answer, which requires more strong reasoning ability. Therefore, we assign a weight of $5$ to each ES question and $10$ to each CA question. Human evaluations are often necessary to accurately assess the quality of answers to these questions but incur huge labor costs. Alternatively, we introduce an automated scoring scheme based on LLMs (Zhang et al., 2023b; Bai et al., 2023b). We set up prompts for GPT-4o to act as a referee to rate the responses of all other LLMs. Furthermore, we meticulously assign a score to every sub-question within a CA question, enabling the referee to furnish precise rationales for their grading decisions. We also manually score the responses of some LLMs for CA questions and compute the Spearman and Kendall correlation coefficients to quantify the agreement between the human and GPT-4o scoring results in Table 3, which indicates that the scores provided by GPT-4o are mostly consistent with those of human evaluations.
@@
Domain Model ToF MC PA ES CA $S_{\text{total}}$ $S_{\text{total}}$ (Scaled)
Medical LLM BianQue-7B 0 0 0 3394 150 3544 25.32
QiZhenGPT-13B 0 0 0 4141.3 170 4311.3 30.80
PULSE-7B 139 1021 95 4009 213 5477 39.13
PULSE-20B 151 1695.5 145 3977.4 248 6216.9 44.42
Open-Source LLM Baichuan2-7B 156.5 1641.5 108 4373.3 284.5 6563.8 46.90
Baichuan2-13B 158 1771.5 132 4814.7 312 7188.2 51.36
ChatGLM3-6B 146.5 1312.5 120 4199.7 231 6009.7 42.94
InternLM2-7B 153.5 1782 188 5257.4 351 7731.9 55.25
InternLM2-20B 140 1839.5 213 5345.3 384 7921.8 56.60
LLaMa3-8B 140 1322.5 142 3565 148 5317.5 37.99
LLaMa3-70B 196 2398.5 404 4640.7 287 7926.2 56.63
Qwen1.5-7B 158.5 1392.5 143 3957.2 252 5903.2 42.18
Qwen1.5-14B 191 2190 376 4335.8 326 7418.8 53.01
Qwen-72B 203.5 3349.5 480 4964 425 9422 67.32
Mixtral-8x7B (47B) 148 1387 181 3834.5 178 5728.5 40.93
Mixtral-8x22B (141B) 177.5 1674 335 4362.9 250 6799.4 48.58
Commercial LLM GLM-4 209.5 3059.5 458 5901.8 443 10071.8 71.96
ERNIE-3.5-8K-0329 230 2949 652 5671.2 514 10016.2 71.57
Qwen-MAX 227.5 3280.5 609 6036.5 515 10668.5 76.23
GPT3.5-turbo 164 1644 324 4563.4 300.5 6995.9 49.99
Full Score 300 4351.5 849 7825 670 13995.5 100
Table 4: Results of different LLMs for the scores for five question types and the total scores on PediaBench. Here, the highest score of each question type is highlighted in bold.
@@ 4.1 Experimental Setup
We validate PediaBench through experiments with 20 general-purpose and medical LLMs, including open-source and commercial models of various scales. Specifically, the models we use include: (1) medical LLMs such as BianQue (Chen et al., 2023a), QiZhenGPT CMKRG et al. (2024), PULSE-7B, and PULSE-20B (Zhang et al., 2023a); (2) the GLM series (Zeng et al., 2023), including ChatGLM3-6B and GLM4; (3) the Baichuan series (Yang et al., 2023), comprising Baichuan2-7B and Baichuan2-13B; (4) the Qwen series (Bai et al., 2023a), including Qwen1.5-7B, Qwen1.5-14B, Qwen1.5-72B, and Qwen-MAX; (5) the InternLM series (Cai et al., 2024b), with InternLM2-7B and InternLM2-20B; (6) the LLaMA3 series (Touvron et al., 2023), featuring LLaMA3-8B and LLaMA3-70B; (7) the Sparse Mixture of Experts language models, namely Mixtral 8x7B and Mixtral 8x22B (Jiang et al., 2024); (8) GPT3.5-turbo; and (9) ERNIE-3.5-8K-0329 (Sun et al., 2021). We deploy and evaluate open-source LLMs locally using pretrained weights on a server with eight NVIDIA RTX A6000 GPUs. We access and evaluate commercial LLMs through their official APIs. More detailed information on these LLMs is given in Appendix B.
@@ 4.2 Evaluation Results
The results for the overall performance of LLMs are shown in Table 4. We present the measures of each LLM for the five types of questions. Note that GPT-4o can achieve the highest scores for three types of objective questions but is omitted from Table 4 to avoid unfair comparison since GPT-4o serves as a referee to assess subjective questions. BianQue-7B and QiZhenGPT-13B cannot correctly understand and follow the instructions for the objective questions, and thus, their scores are 00 for these types. In terms of ToF, PULSE-7B obtains the lowest score of 139 among all models, which is not even half of the total score for ToF. ERNIE-3.5-8K-0329 generally has the best performance for ToF and PA. Generally, larger models achieve better performance than smaller ones in both objective and subjective questions. However, GPT3.5-turbo exhibited a notably subpar performance, obtaining only half the full score. Such results are quite surprising considering the strong capacities of GPT3.5-turbo in general and medical linguistic tasks. We conjecture that this is because the GPT3.5-turbo training data lacks language resources in Chinese.
(a) Scores (scaled to 0–100) of LLMs for five question types
@@
  • Until now, only several commercial LLMs and Qwen-72B have achieved passing overall scores (i.e., $\geq 60$). Qwen-MAX achieved the highest total score of 76.23, firmly securing its top position but still far from achieving excellence in the exam (note that a total score of GPT-4o is nearly 80, which is not presented due to a potentially unfair comparison). Medical LLMs received lower scores due to poor instruction-following, reasoning, and generation capacity. On the one hand, this confirms that PediaBench is challenging and discriminative enough to assess the applicability of LLMs in pediatric scenarios. On the other hand, this also indicates that, despite the commendable efforts, there remains a huge gap to meet the requirement for deploying LLMs as AI assistants for pediatricians.
  • Small LLMs can sometimes show good performance in the PediaBench dataset. For example, Qwen1.5-14B, with only 14B parameters, outperforms GPT3.5-turbo and Mixtral-8x22B (141B) with many more parameters. This could potentially be explained by the fact that GPT3.5-turbo and Mixtral-8x22B (141B) may lack sufficient exposure to Chinese medical corpora during their training. However, large-scale and commercial models exhibit significantly better performance in most cases.
@@ 5 Conclusion
In this paper, we introduce PediaBench, a comprehensive Chinese benchmark dataset encompassing 12 pediatric disease groups to evaluate the QA capacity of LLMs. Specifically, PediaBench consists of 4,565 objective and 1,632 subjective questions. It adopts an integrated scoring criterion based on different difficulty levels to thoroughly assess the proficiency of an LLM in pediatric problems. Finally, the effectiveness of PediaBench is confirmed through extensive experiments on 20 open-source and commercial LLMs. Through an in-depth analysis of experimental results, we offer insights into the ability of LLMs for pediatric QA in Chinese, highlighting their limitations for improvement.
@@ Appendix D Experimental Results for Different Disease Groups
Table 7 shows the scores of each LLM specifically for objective questions and ES questions across 12 disease groups. In order to quantify the score, we calculated the proportion of LLMs’ score based on the inconsistent number of questions and scores for each disease group. Due to the inadequate instruction-following capabilities of BianQue (7B) and QiZhenGPT (13B), they struggled to provide precise responses to objective questions, resulting in surprisingly low scores. Intriguingly, despite Mixtral-8x7B’s (47B) commendable instruction-following abilities, its error rate in providing answers remained remarkably high, ultimately leading to a comparable score with that of BianQue (7B) and QiZhenGPT (13B). We observe that most models achieve their highest scores in the two disease groups of HCDA of DImS. Questions related to HCDA often involve basic medical knowledge, which does not require in-depth specialization. Moreover, many diseases in the DImS group share similar clinical manifestations and treatment strategies, providing models with clearer clues for answers. No models can perform well for subjective questions in all performance measures across different groups of diseases.
@@ E.1 True-or-False Question
For true-or-false questions, we compare the answers of PULSE-7B, LLaMA3-8B, Qwen-7B-Chat, and ERNIE-3.5-8K-0329, as shown in Figure 8. After in-depth analysis, we found an interesting phenomenon: models with smaller parameters, such as Qwen-7B-Chat, could judge statements correctly, even though there are errors in their explanations. Furthermore, despite our restrictions on the language generated by the models, the output of LLaMA3-8B always contains a mixture of Chinese and English.
@@ E.2 Multi-Choice Question
We conducted an in-depth analysis of a multiple-choice question on the disease of the infectious system. The answers provided by several LLMs are shown in Figure  9. In this case, LLaMA3-8B and PULSE-20B provided incorrect options and explanations, failing to capture the key points of the question accurately. Baichuan2-7B was even unable to make a reasonable choice. However, ERNIE-3.5-8K-0329 not only selected the correct answer but also provided a detailed and reasonable explanation.

### v3

@@ Abstract
The emergence of Large Language Models (LLMs) in the medical domain has stressed a compelling need for standard datasets to evaluate their question-answering (QA) performance. Although there have been several benchmark datasets for medical QA, they either cover common knowledge across different departments or are specific to another department rather than pediatrics. Moreover, some of them are limited to objective questions and do not measure the generation capacity of LLMs. Therefore, they cannot comprehensively assess the QA ability of LLMs in pediatrics. To fill this gap, we construct PediaBench, the first Chinese pediatric dataset for LLM evaluation. Specifically, it contains 4,117 objective questions and 1,632 subjective questions spanning 12 pediatric disease groups. It adopts an integrated scoring criterion based on different difficulty levels to thoroughly assess the proficiency of an LLM in instruction following, knowledge understanding, clinical case analysis, etc. Finally, we validate the effectiveness of PediaBench with extensive experiments on 20 open-source and commercial LLMs. Through an in-depth analysis of experimental results, we offer insights into the ability of LLMs to answer pediatric questions in the Chinese context, highlighting their limitations for further improvements. Our code and data are published anonymously at https://github.com/ACMISLab/PediaBench.
@@ 1 Introduction
To address the above challenges, in this paper, we introduce PediaBench, the first comprehensive Chinese benchmark dataset in pediatrics. As shown in Figure 1, PediaBench contains 4,117 objective questions and 1,632 subjective questions collected from various sources, such as the Chinese National Medical Licensing Examination (CNMLE), final exams of universities in medicine, pediatric disease diagnosis and treatment standards, and clinical guidelines. PediaBench encompasses five distinct types of questions, namely true or false, multiple choice, pairing, essay/short answer, and case analysis, across 12 typical pediatric disease groups. Furthermore, to provide an accurate evaluation of the performance of each LLM for QA in pediatrics, we use a scoring criterion that combines difficulty levels and automatic scoring. On the one hand, we assign different scores to objective questions based on the difficulty factor; on the other hand, we adopt the Language-Model-as-an-Examiner (LME) framework Bai et al. (2023b) to score the objective questions based on GPT-4o. Finally, we conduct extensive experiments with 20 LLMs on the PediaBench dataset and provide an in-depth analysis of the performance of LLMs for QA in pediatrics. Our main contributions are summarized as follows:
@@ 3.4 Evaluation Criteria
For ToF and MC, we use accuracy as a basic performance measure for each question. Then, we design a scoring scheme based on difficulty coefficient to assess the overall performance of each LLM. We calculate the difficulty coefficient $Dc_{i}$ of each question $i$ based on the accuracy of all LLMs’ answers. Specifically, we have $Dc_{i}=1-\frac{\text{num}^{\text{acc}}_{i}}{\text{num}^{\text{llm}}_{i}}$, where $\text{num}^{\text{acc}}_{i}$ is the number of LLMs that correctly answer question $i$ and $\text{num}^{\text{llm}}_{i}$ is the total number of LLMs in the evaluation. All questions are then divided into four difficulty levels, namely simple, normal, difficult, and extreme, with different weights $w_{i}$ based on $Dc_{i}$ in the scoring scheme as follows:
@@
Figure 3 presents the number of questions at each difficulty level in ToF and MC. For PA, we assign a weight of $w_{i}=3$ to each question $i$ and use the following scoring rules: (1) A completely correct answer gets $3$ points; (2) a partially correct answer gets $1$ point; and (3) a completely incorrect answer does not get any points.
@@
ES and CA are open-ended questions with no unique answers. CA is even harder than ES because it may require the integration of multiple pieces of knowledge to answer, which requires more strong reasoning ability. Therefore, we assign a weight of $5$ to each ES question and $10$ to each CA question. Human evaluations are often necessary to accurately assess the quality of answers to these questions but incur huge labor costs. Alternatively, we introduce an automated scoring scheme based on LLMs (Zhang et al., 2023b; Bai et al., 2023b). We set up prompts for GPT-4o to act as a referee to rate the responses of all other LLMs. Furthermore, we meticulously assign a score to every sub-question within a CA question, enabling the referee to furnish precise rationales for their grading decisions. We also manually score the responses of some LLMs for CA questions and compute the Spearman and Kendall correlation coefficients to quantify the agreement between the human and GPT-4o scoring results in Table 3, which indicates that the scores provided by GPT-4o are mostly consistent with those of human evaluations.
Finally, the total score of an LLM for each type of question is the sum of the scores of all (partially) correctly answered questions. We follow the common practice of standardized medical examinations to assign a fixed proportion to each type of question (10 for ToF, 40 for MC, 10 for PA, 30 for ES, and 10 for CA). The integrated score of an LLM is thus calculated as the weighted sum of the total score for each question type, i.e., $\mathrm{S}_{\text{total}}=\mathrm{S}_{\text{ToF}}\times{10}+\mathrm{S}_{\text{MC}}\times{40}+\mathrm{S}_{\text{PA}}\times{10}+\mathrm{S}_{\text{ES}}\times{30}+\mathrm{S}_{\text{CA}}\times{10}$.
@@
Domain Model ToF MC PA ES CA $\mathrm{S}_{\text{total}}$
Medical LLM BianQue (7B) 0 0 0 13.01 2.24 15.25
QiZhenGPT (13B) 0 0 0 15.88 2.54 18.41
PULSE (7B) 4.63 9.39 1.12 15.37 3.18 33.69
PULSE (20B) 5.03 15.59 1.71 15.25 3.70 41.28
Open-Source LLM Baichuan2-7B 5.22 15.09 1.27 16.77 4.25 42.59
Baichuan2-13B 5.27 16.28 1.55 18.46 4.66 46.22
ChatGLM3-6B 4.88 12.06 1.41 16.10 3.45 37.91
InternLM2-7B 5.12 16.38 2.21 20.16 5.24 49.11
InternLM2-20B 4.67 16.91 2.51 20.49 5.73 50.31
LLaMa3-8B 4.67 12.16 1.67 13.67 2.21 34.37
LLaMa3-70B 6.53 22.05 4.76 17.79 4.28 55.41
Qwen1.5-7B 5.28 12.80 1.68 15.17 3.76 38.70
Qwen1.5-14B 6.37 20.13 4.43 16.62 4.87 52.41
Qwen-72B 6.78 30.79 5.65 19.03 6.34 68.60
Mixtral-8x7B 4.93 12.75 2.13 14.70 2.66 37.17
Mixtral-8x22B 5.92 15.39 3.95 16.73 3.73 45.71
Commercial LLM GLM-4 6.98 28.12 5.39 22.63 6.61 69.74
ERNIE-3.5 7.67 27.11 7.68 21.74 7.67 71.87
Qwen-MAX 7.58 30.16 7.17 23.14 7.69 75.74
GPT3.5-turbo 5.47 15.11 3.82 17.50 4.49 46.38
Full Score 10 40 10 30 10 100
Table 4: Scores of different LLMs for the five question types on the PediaBench dataset.
@@ 4.1 Experimental Setup
We validate PediaBench through experiments with 20 general-purpose and medical LLMs, including open-source and commercial models of various scales. Specifically, the models we use include: (1) medical LLMs such as BianQue (Chen et al., 2023a), QiZhenGPT CMKRG et al. (2024), PULSE-7B, and PULSE-20B (Zhang et al., 2023a); (2) the GLM series (Zeng et al., 2023), including ChatGLM3-6B and GLM4; (3) the Baichuan series (Yang et al., 2023), comprising Baichuan2-7B and Baichuan2-13B; (4) the Qwen series (Bai et al., 2023a), including Qwen1.5-7B, Qwen1.5-14B, Qwen1.5-72B, and Qwen-MAX; (5) the InternLM series (Cai et al., 2024b), with InternLM2-7B and InternLM2-20B; (6) the LLaMA3 series (Touvron et al., 2023), featuring LLaMA3-8B and LLaMA3-70B; (7) the Sparse Mixture of Experts language models, namely Mixtral 8x7B and Mixtral 8x22B (Jiang et al., 2024); (8) GPT3.5-turbo; and (9) ERNIE-3.5 (We choose ERNIE-3.5-8K-0329) (Sun et al., 2021). We deploy and evaluate open-source LLMs locally using pretrained weights on a server with eight NVIDIA RTX A6000 GPUs. We access and evaluate commercial LLMs through their official APIs. More detailed information on these LLMs is given in Appendix B.
@@ 4.2 Evaluation Results
The results for the overall performance of LLMs are shown in Table 4. We present the measures of each LLM for the five types of questions. Note that GPT-4o can achieve the highest scores for three types of objective questions but is omitted from Table 4 to avoid unfair comparison since GPT-4o serves as a referee to assess subjective questions. BianQue-7B and QiZhenGPT-13B cannot correctly understand and follow the instructions for the objective questions, and thus, their scores are 00 for these types. In terms of ToF, PULSE-7B obtains the lowest score of 4.63 among all models, which is not even half of the total score for ToF. ERNIE-3.5 generally has the best performance for ToF and PA. Generally, larger models achieve better performance than smaller ones in both objective and subjective questions. However, GPT3.5-turbo exhibited a notably subpar performance, obtaining only half the full score. Such results are quite surprising considering the strong capacities of GPT3.5-turbo in general and medical linguistic tasks. We conjecture that this is because the GPT3.5-turbo training data lacks language resources in Chinese.
(a) Scores (scaled to 0–100) of open-source LLMs for five question types
@@
  • Until now, only several commercial LLMs and Qwen-72B have achieved passing overall scores (i.e., $\geq 60$). Qwen-MAX achieved the highest total score of 75.74, firmly securing its top position but still far from achieving excellence in the exam (note that a total score of GPT-4o is nearly 80, which is not presented due to a potentially unfair comparison). Medical LLMs received lower scores due to poor instruction-following, reasoning, and generation capacity. On the one hand, this confirms that PediaBench is challenging and discriminative enough to assess the applicability of LLMs in pediatric scenarios. On the other hand, this also indicates that, despite the commendable efforts, there remains a huge gap to meet the requirement for deploying LLMs as AI assistants for pediatricians.
  • Small LLMs can sometimes show good performance in the PediaBench dataset. For example, Qwen1.5-14B, with only 14B parameters, outperforms GPT3.5-turbo and Mixtral-8x22B with many more parameters. This could potentially be explained by the fact that GPT3.5-turbo and Mixtral-8x22B may lack sufficient exposure to Chinese medical corpora during their training. However, large-scale and commercial models exhibit significantly better performance in most cases.
@@ 5 Conclusion
In this paper, we introduce PediaBench, a comprehensive Chinese benchmark dataset encompassing 12 pediatric disease groups to evaluate the QA capacity of LLMs. Specifically, PediaBench consists of 4,117 objective and 1,632 subjective questions. It adopts an integrated scoring criterion based on different difficulty levels to thoroughly assess the proficiency of an LLM in pediatric problems. Finally, the effectiveness of PediaBench is confirmed through extensive experiments on 20 open-source and commercial LLMs. Through an in-depth analysis of experimental results, we offer insights into the ability of LLMs for pediatric QA in Chinese, highlighting their limitations for improvement.
@@ Appendix D Experimental Results for Different Disease Groups
Table 7 shows the scores of each LLM specifically for objective questions and ES questions across 12 disease groups. In order to quantify the score, we calculated the proportion of LLMs’ score based on the inconsistent number of questions and scores for each disease group. Due to the inadequate instruction-following capabilities of BianQue (7B) and QiZhenGPT (13B), they struggled to provide precise responses to objective questions, resulting in surprisingly low scores. Intriguingly, despite Mixtral-8x7B’s commendable instruction-following abilities, its error rate in providing answers remained remarkably high, ultimately leading to a comparable score with that of BianQue (7B) and QiZhenGPT (13B). We observe that most models achieve their highest scores in the two disease groups of HCDA of DImS. Questions related to HCDA often involve basic medical knowledge, which does not require in-depth specialization. Moreover, many diseases in the DImS group share similar clinical manifestations and treatment strategies, providing models with clearer clues for answers. No models can perform well for subjective questions in all performance measures across different groups of diseases.
@@ E.1 True-or-False Question
For true-or-false questions, we compare the answers of PULSE-7B, LLaMA3-8B, Qwen-7B-Chat, and ERNIE-3.5, as shown in Figure 8. After in-depth analysis, we found an interesting phenomenon: models with smaller parameters, such as Qwen-7B-Chat, could judge statements correctly, even though there are errors in their explanations. Furthermore, despite our restrictions on the language generated by the models, the output of LLaMA3-8B always contains a mixture of Chinese and English.
@@ E.2 Multi-Choice Question
We conducted an in-depth analysis of a multiple-choice question on the disease of the infectious system. The answers provided by several LLMs are shown in Figure  9. In this case, LLaMA3-8B and PULSE-20B provided incorrect options and explanations, failing to capture the key points of the question accurately. Baichuan2-7B was even unable to make a reasonable choice. However, ERNIE-3.5 not only selected the correct answer but also provided a detailed and reasonable explanation.